# It’s a (Blind) Match! Towards Vision-Language Correspondence without Parallel Data

### v1

@@ 5.1 Small-scale matching
Results. We show the matching accuracy for different vision models in Fig. 4. On the left, we use all-mpnet-base-v2 evaluated on CIFAR-10 and, on the right, All-Roberta-large-v1 on CINIC-10. More combinations are given in Sec. F.1. We observe that most models perform better than 10% accuracy, which is the baseline for random permutations. Moreover, the DINOv2 model can be matched with the highest accuracy of 72% on CIFAR-10 and 100% on CINIC-10. Curiously, we observe that it is the pre-training strategy that seems to have a larger impact than the model size. On average for all language models, the DINOv2 models are 5.2% better on CIFAR-10 and 7.7% better on CINIC-10 than the second-best pre-training strategy.
@@ 5.3 Solver comparison
Solver | Accuracy (%) | Cost | Global? (%) |
Random | ±10 | ±0.30 | ±0 |
LocalCKA [30] | ±9 | ±0.27 | ±0 |
OT [38] | ±4 | ±0.11 | ±0 |
FAQ [49] | ±3 | ±0.08 | ±0 |
MPOpt [18] | ±12 | ±0.02 | ±51 |
Gurobi [12] | ±0 | ±0.02 | ±0 |
Ours | ±0 | ±0.02 | ±0 |
Results. In the small-scale setting, we can compute the global optimum with enumeration. In Tab. 1, we compare the matching accuracy, the cost after optimization, and the percentage that the global optimum is reached for every method. For LocalCKA, the CKA metric is minimized using a linear approximation. We observe that minimizing this approximation does not lead to better results than random permutations, both in terms of Gromov-Wasserstein loss and the matching accuracy. Both local solvers (OT, FAQ) are able to minimize the loss, but are stuck in a local minimum, thus their accuracy is not better than the random baseline. This shows that local optima are not sufficient to find a useful match. MPOpt manages to approach the optimal objective, but only reaches it about half the time. In contrast, both our heuristic and Gurobi find the global optimum every time, resulting in accuracy of 60%.
Fig. 6 reports the results on the larger-scale problem. We show the cost and the lower bound for each problem size. For ease of comparison and for computational reasons, we select only one set of classes for each problem size. Consistent with the small-scale experiment, LocalCKA does not perform better than random and produces out-of-bound values. Interestingly, OT and FAQ are able to solve some instances with global optimality. We suspect that this is because the optimal problem may have a specialized structure. Therefore, the relaxations can lead to the global optimum for such problems. Already at size , all but our method fail to reach the global optimum. For problem sizes larger than 60, Gurobi does not produce a solution within the given time limit. Our method consistently provides the tightest bounds, provably leading to a global optimum up to size . Furthermore, it achieves the best primal solution for most sizes, even surpassing Gurobi on problems of size .
@@ 5.4 Application: Unsupervised classifiers
Model | All-Roberta-large-v1 (%) | all-mpnet-base-v2 (%) |
CLIP | ±0.5 | ±0.4 |
DeiT | ±3.4 | ±7.5 |
DINOv2 | ±9.2 | ±0.4 |
Results. We evaluate our unsupervised classifier with several vision and language models in Tab. 2. The best accuracy is achieved by the DINOv2 model with 51.5% for All-Roberta-large-v1. However, all of the considered models achieve a higher accuracy than the random baseline of 10%. These results are clearly inferior to the supervised models. Nevertheless, to the best of our knowledge, this is the first instance of fully unsupervised classification. Furthermore, we show in Sec. F.2 that previous solvers are not able to find meaningful matches in this setting and analyze how clustering changes the matching performance.
@@ Appendix B Mutual k-NN and CKA as distortion metrics
Models | CIFAR-10 (%) | CINIC-10 (%) |
Mutual k-NN | ||
CKA | ||
@@
Tab. 3 empirically compares these formulations in a small-scale data regime, using CIFAR-10 [23] and CINIC-10 [9] datasets introduced in the main text. We find that despite the wider adoption of Mutual k-NN and CKA metrics in previous work, the Gromov-Wasserstein distance leads to a higher matching accuracy on both datasets.
@@ D.2 Implementation details
Hahn-Grant | -1.979877 | – | ||
+ factorized | -1.979875 | -2.089994 | ||
+ auction | -1.979871 | -1.981942 | ||
+ Jonker-Volgenant | -1.979878 | -1.980387 | ||
+ LAP solutions | -1.979914 | -1.980387 | ||
+ primal heuristic | -1.980015 | -1.980387 |
@@ D.3 Ablations
Results. Tab. 4 shows the result for each adaptation for size and Tab. 5 for size . We observe that the original Hahn-Grant solver reaches a non-trivial primal-dual bound for size but does not finish the first iteration within the two hour time limit for size . Even though factorization was introduced for improved memory, it also slightly speeds up the computation, leading to a smaller bound for both sizes. The small increase in the objective for size can be explained by the fact that the returned primal estimate is only the primal solution of the LAP for the leader. Therefore, the quality of this estimate can vary with every iteration. The auction and Jonker-Volgenant algorithms were introduced to improve the speed of the algorithm and, therefore, lead to smaller bounds. However, the Jonker-Volgenant algorithm leads to the tightest bounds empirically for large problems while the auction algorithm is slightly better on the size problem. The LAP solutions do not change the dual estimate but improve the primal solution by a significant margin. Finally, the primal heuristics further improve the primal solution. Since we are mostly interested in good primal solutions and measuring the quality of the solution, we use the algorithm with all the adaptations.
@@
Hahn-Grant | -1.950160 | -2.245823 | ||
+ factorized | -1.950166 | -1.951418 | ||
+ auction | -1.950197 | -1.950586 | ||
+ Jonker-Volgenant | -1.950197 | -1.950623 | ||
+ LAP solutions | -1.950268 | -1.950623 | ||
+ primal heuristic | -1.950360 | -1.950623 |
@@ Appendix E Experimental details
Solver | Accuracy (%) | Matching Accuracy (%) |
Random | ±10.3 | ±11.6 |
LocalCKA | ±7.0 | ±7.7 |
OT | ±9.5 | ±9.4 |
FAQ | ±8.7 | ±9.9 |
MPOpt | ±21.4 | ±23.9 |
Ours | ±3.4 | ±3.7 |
GT | ±0.6 | ±0.0 |
@@ F.1 Comparison of vision and language models
In Fig. 11, we show the matching accuracy (top) and Mutual k-NN with (bottom) of each vision model in combination with All-Roberta-large-v1 (left) and all-mpnet-base-v2 (right) for CIFAR-10 (left) and CINIC-10 (right). The lines show the trend for increasing model sizes that are fitted to the models of varying capacity for each model class (different colors). The model size corresponds to the number of parameters. First, we observe that for both datasets and language models, DINOv2 yields the highest matching accuracy. Furthermore, for every model class, there is no a clear propensity of larger models to perform better. As this seems to contradict the platonic representation hypothesis Huh et al. [17], we report the Mutual k-NN in the bottom plots. We observe that in line with our conclusions, there is still no clear improvement of the Mutual k-NN metric w.r.t. an increasing model capacity. This implies that larger models do not necessarily lead to stronger alignment with language – at least on CIFAR-10 and CINIC-10. Considering the observations by Huh et al. [17], this suggests that scaling the models could lead to a better alignment on the Wit dataset [6], even though it may not be sufficient to improve the alignment on every other dataset. We report the individual matching accuracy of each combination of vision and language model for CIFAR-10 in Fig. 12 and for CINIC-10 in Fig. 13.
@@ F.3 Comparison of solvers on small-scale problems
Similar to Tab. 1, we benchmark different solvers for all combinations in Tab. 7. Our factorized Hahn-Grant solver and Gurobi [12] find the global optimum for all problems. For most combinations, these solvers are also the only solvers finding the global optimum. MPOpt [18] also finds the global optimum for some problems (e.g., CIFAR-10 with DINOv2) but fails for other problems (e.g., CIFAR-10 with CLIP). For all but two experiments (CIFAR-10 with DeiT and all-mpnet-base-v2 and CINIC-10 with DeiT and All-Roberta-large-v1), the global optimum also leads to the best matching accuracy, which underlines the importance of the Gromov-Wasserstein distance as a proxy measure for blind matching. Finally, the matching accuracy is close to random for most solvers (LocalCKA [30], OT [38], and FAQ [49]). This shows that a global solver is crucial to obtain non-trivial results.
All-Roberta-large-v1 | all-mpnet-base-v2 | |||||||
Accuracy (%) | Cost | Global? (%) | Accuracy (%) | Cost | Global? (%) | |||
CIFAR-10 | CLIP | Random | ±10.4 | ±0.32 | ±0.0 | ±10.4 | ±0.29 | ±0.0 |
LocalCKA [30] | ±8.87 | ±0.33 | ±0.0 | ±11.64 | ±0.31 | ±0.0 | ||
OT [38] | ±3.08 | ±0.03 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
FAQ [49] | ±3.66 | ±0.06 | ±0.0 | ±8.89 | ±0.02 | ±41.04 | ||
MPOpt [18] | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±10.05 | ±0.01 | ±0.0 | ±4.44 | ±0.01 | ±0.0 | ||
Ours | ±10.05 | ±0.01 | ±0.0 | ±4.44 | ±0.01 | ±0.0 | ||
DeiT | Random | ±10.4 | ±0.29 | ±0.0 | ±10.4 | ±0.27 | ±0.0 | |
LocalCKA [30] | ±9.18 | ±0.2 | ±0.0 | ±6.81 | ±0.22 | ±0.0 | ||
OT [38] | ±0.0 | ±0.01 | ±0.0 | ±3.08 | ±0.01 | ±47.02 | ||
FAQ [49] | ±10.7 | ±0.13 | ±0.0 | ±14.9 | ±0.01 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.02 | ±0.0 | ±0.0 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±4.7 | ±0.01 | ±0.0 | ±6.81 | ±0.0 | ±0.0 | ||
Ours | ±4.7 | ±0.01 | ±0.0 | ±6.81 | ±0.0 | ±0.0 | ||
DINOv2 | Random | ±10.4 | ±0.27 | ±0.0 | ±10.4 | ±0.28 | ±0.0 | |
LocalCKA [30] | ±11.42 | ±0.31 | ±0.0 | ±9.1 | ±0.24 | ±0.0 | ||
OT [38] | ±26.25 | ±0.44 | ±0.0 | ±24.15 | ±0.54 | ±22.36 | ||
FAQ [49] | ±35.3 | ±0.18 | ±22.36 | ±23.82 | ±0.22 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
Ours | ±0.0 | ±0.01 | ±0.0 | ±17.65 | ±0.01 | ±0.0 | ||
CINIC-10 | CLIP | Random | ±10.4 | ±0.31 | ±0.0 | ±10.4 | ±0.3 | ±0.0 |
LocalCKA [30] | ±8.87 | ±0.25 | ±0.0 | ±8.58 | ±0.27 | ±0.0 | ||
OT [38] | ±3.66 | ±0.13 | ±0.0 | ±3.66 | ±0.11 | ±0.0 | ||
FAQ [49] | ±5.5 | ±0.02 | ±0.0 | ±3.08 | ±0.08 | ±0.0 | ||
MPOpt [18] | ±2.24 | ±0.04 | ±22.36 | ±12.1 | ±0.02 | ±51.3 | ||
Gurobi [12] | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.02 | ±0.0 | ||
Ours | ±0.0 | ±0.01 | ±0.0 | ±0.0 | ±0.02 | ±0.0 | ||
DeiT | Random | ±10.4 | ±0.25 | ±0.0 | ±10.4 | ±0.25 | ±0.0 | |
LocalCKA [30] | ±7.88 | ±0.17 | ±0.0 | ±8.75 | ±0.22 | ±0.0 | ||
OT [38] | ±6.57 | ±0.02 | ±0.0 | ±4.7 | ±0.05 | ±0.0 | ||
FAQ [49] | ±17.01 | ±0.07 | ±0.0 | ±3.08 | ±0.04 | ±0.0 | ||
MPOpt [18] | ±4.89 | ±0.01 | ±51.04 | ±6.16 | ±0.01 | ±0.0 | ||
Gurobi [12] | ±3.08 | ±0.01 | ±0.0 | ±6.16 | ±0.01 | ±0.0 | ||
Ours | ±3.08 | ±0.01 | ±0.0 | ±6.16 | ±0.01 | ±0.0 | ||
DINOv2 | Random | ±10.4 | ±0.23 | ±0.0 | ±10.4 | ±0.23 | ±0.0 | |
LocalCKA [30] | ±9.45 | ±0.14 | ±0.0 | ±7.33 | ±0.21 | ±0.0 | ||
OT [38] | ±15.72 | ±0.06 | ±48.94 | ±5.03 | ±0.22 | ±0.0 | ||
FAQ [49] | ±13.33 | ±0.06 | ±0.0 | ±10.21 | ±0.12 | ±0.0 | ||
MPOpt [18] | ±0.0 | ±0.02 | ±22.36 | ±13.8 | ±0.03 | ±30.78 | ||
Gurobi [12] | ±0.0 | ±0.02 | ±0.0 | ±7.33 | ±0.02 | ±0.0 | ||
Ours | ±0.0 | ±0.02 | ±0.0 | ±7.33 | ±0.02 | ±0.0 |

### v2

@@ 5.1 Small-scale matching
Results. We show the matching accuracy for different vision models in Fig. 4. On the left, we use all-mpnet-base-v2 evaluated on CIFAR-10 and, on the right, All-Roberta-large-v1 on CINIC-10. More combinations are given in Sec. F.1. We observe that most models perform better than 10% accuracy, which is the baseline for random permutations. Moreover, the DINOv2 model can be matched with the highest accuracy of 80% on CIFAR-10 and 100% on CINIC-10. Curiously, we observe that it is the pre-training strategy that seems to have a larger impact than the model size. On average for all language models, the DINOv2 models are 5.3% better on CIFAR-10 and 7.6% better on CINIC-10 than the second-best pre-training strategy.
@@ 5.3 Solver comparison
Solver | Accuracy (%) | Cost | Global? (%) | |||
Random | ||||||
LocalCKA [30] | ||||||
OT [38] | ||||||
FAQ [49] | ||||||
MPOpt [18] | ||||||
Gurobi [12] | ||||||
Ours |
Results. In the small-scale setting, we can compute the global optimum with enumeration. In Tab. 1, we compare the matching accuracy, the cost after optimization, and the percentage that the global optimum is reached for every method. For LocalCKA, the CKA metric is minimized using a linear approximation. We observe that minimizing this approximation does not lead to significant better results than random permutations in terms of the matching accuracy. Both local solvers (OT, FAQ) are able to minimize the loss, but are stuck in a local minimum, thus their accuracy is often not much better than the random baseline. This shows that local optima are not sufficient to find a useful match. MPOpt manages to approach the optimal objective, but only reaches it about 90% of the time. In contrast, both our heuristic and Gurobi find the global optimum every time, resulting in an accuracy of 100%.
Fig. 6 reports the results on the larger-scale problem. We show the cost and the lower bound for each problem size. For ease of comparison and for computational reasons, we select only one set of classes for each problem size. Consistent with the small-scale experiment, LocalCKA produces the worst values. Interestingly, OT and FAQ are able to solve some instances with global optimality. We suspect that this is because the optimal problem may have a specialized structure. Therefore, the relaxations can lead to the global optimum for such problems. Already at size , all but our method fail to reach the global optimum. For problem sizes larger than 40, Gurobi does not produce a solution within the given time limit. Our method consistently provides the tightest bounds, provably leading to a global optimum up to size . Furthermore, it achieves the best primal solution for most sizes.
@@ 5.4 Application: Unsupervised classifiers
Model | All-Roberta-large-v1 (%) | all-mpnet-base-v2 (%) | ||
CLIP | ||||
DeiT | ||||
DINOv2 |
Results. We evaluate our unsupervised classifier with several vision and language models in Tab. 2. The best accuracy is achieved by the DINOv2 model with 51.1% for All-Roberta-large-v1. However, all of the considered models achieve a higher accuracy than the random baseline of 10%. These results are clearly inferior to the supervised models. Nevertheless, to the best of our knowledge, this is the first instance of fully unsupervised classification. Furthermore, we show in Sec. F.2 that previous solvers are not able to find meaningful matches in this setting and analyze how clustering changes the matching performance.
@@ Appendix B Mutual k-NN and CKA as distortion metrics
Models | CIFAR-10 (%) | CINIC-10 (%) | ||
Mutual k-NN | ||||
CKA | ||||
@@
Tab. 3 empirically compares these formulations in a small-scale data regime, using CIFAR-10 [23] and CINIC-10 [9] datasets introduced in the main text. We find that despite the wider adoption of Mutual k-NN and CKA metrics in previous work, the Gromov-Wasserstein distance leads to a comparable or higher matching accuracy on both datasets.
@@ D.2 Implementation details
Hahn-Grant | – | |||
+ factorized | ||||
+ auction | ||||
+ Jonker-Volgenant | ||||
+ LAP solutions | ||||
+ primal heuristic |
@@ D.3 Ablations
Results. Tab. 4 shows the result for each adaptation for size and Tab. 5 for size . We observe that the original Hahn-Grant solver reaches a non-trivial primal-dual bound for size but does not finish the first iteration within the two hour time limit for size . Even though factorization was introduced for improved memory, it also slightly speeds up the computation, leading to a smaller bound for both sizes. The auction and Jonker-Volgenant algorithms were introduced to improve the speed of the algorithm and, therefore, lead to smaller bounds. However, the Jonker-Volgenant algorithm leads to the tightest bounds empirically for large problems while the auction algorithm is slightly better on the size problem. The LAP solutions do not change the dual estimate but improve the primal solution by a significant margin. Finally, the primal heuristics further improve the primal solution with the trade-off of slightly worse bounds. Since we are mostly interested in good primal solutions and measuring the quality of the solution, we use the algorithm with all the adaptations.
@@
Hahn-Grant | ||||
+ factorized | ||||
+ auction | ||||
+ Jonker-Volgenant | ||||
+ LAP solutions | ||||
+ primal heuristic |
@@ Appendix E Experimental details
Solver | Accuracy (%) | Matching Accuracy (%) | ||
Random | ||||
LocalCKA [30] | ||||
OT [38] | ||||
FAQ [49] | ||||
MPOpt [18] | ||||
Ours | ||||
GT |
@@ F.1 Comparison of vision and language models
In Fig. 11, we show the matching accuracy (top) and Mutual k-NN with (bottom) of each vision model in combination with All-Roberta-large-v1 (left) and all-mpnet-base-v2 (right) for CIFAR-10 (left) and CINIC-10 (right). The lines show the trend for increasing model sizes that are fitted to the models of varying capacity for each model class (different colors). The model size corresponds to the number of parameters. First, we observe that for both datasets and language models, DINOv2 yields the highest matching accuracy. Furthermore, for every model class, there is no a clear propensity of larger models to perform better. As this seems to contradict the platonic representation hypothesis [17], we report the Mutual k-NN in the bottom plots. We observe that in line with our conclusions, there is still no clear improvement of the Mutual k-NN metric w.r.t. an increasing model capacity. This implies that larger models do not necessarily lead to stronger alignment with language – at least on CIFAR-10 and CINIC-10. Considering the observations by Huh et al. [17], this suggests that scaling the models could lead to a better alignment on the Wit dataset [6], even though it may not be sufficient to improve the alignment on every other dataset. We report the individual matching accuracy of each combination of vision and language model for CIFAR-10 in Fig. 12 and for CINIC-10 in Fig. 13.
@@ F.3 Comparison of solvers on small-scale problems
Similar to Tab. 1, we benchmark different solvers for all combinations in Tab. 7. Our factorized Hahn-Grant solver and Gurobi [12] find the global optimum for all problems. For most combinations, these solvers are also the only solvers finding the global optimum. MPOpt [18] also finds the global optimum for some problems (e.g., CINIC-10 with DINOv2) but fails for other problems (e.g., CIFAR-10 with CLIP). For all but four experiments (all-mpnet-base-v2 on CIFAR-10 for all three vision models and on CINIC-10 with DeiT), the global optimum also leads to the best matching accuracy, which underlines the importance of the Gromov-Wasserstein distance as a proxy measure for blind matching. Finally, the matching accuracy is close to random for most solvers (LocalCKA [30], OT [38], and FAQ [49]). This shows that a global solver is crucial to obtain non-trivial results.
all-mpnet-base-v2 | All-Roberta-large-v1 | |||||||||||||
Accuracy (%) | Cost | Global? (%) | Accuracy (%) | Cost | Global? (%) | |||||||||
CIFAR-10 | CLIP | Random | ||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours | ||||||||||||||
DeiT | Random | |||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours | ||||||||||||||
DINOv2 | Random | |||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours | ||||||||||||||
CINIC-10 | CLIP | Random | ||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours | ||||||||||||||
DeiT | Random | |||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours | ||||||||||||||
DINOv2 | Random | |||||||||||||
LocalCKA [30] | ||||||||||||||
OT [38] | ||||||||||||||
FAQ [49] | ||||||||||||||
MPOpt [18] | ||||||||||||||
Gurobi [12] | ||||||||||||||
Ours |